OpenStudy (anonymous):

simplify the expression and write in standard form (square root of -50)

OpenStudy (anonymous):

-10

OpenStudy (sonic):

\[\sqrt{-50}=\sqrt{-1}*\sqrt{25*2}=i*5\sqrt{2}=5i \sqrt{2}\]